c. 12.5 ml of H3PO4 requires 47.5 ml of 0.10 M KOH. Find molarity for the H3PO4.
1
Expert's answer
2020-06-11T09:23:37-0400

H3PO4 + 3KOH --> K3PO4 + 3H2O

n (KOH) = (47.5mL x 0.10mol) / 1000 mL = 0.00475 moles of KOH present in 47.5 mL of the solution.

According to the equation, n (H3PO4) = n (KOH) / 3.

The molarity is: ((0.00475mol / 3) x 1000mL) / 12.5mL = 0.127M

Answer: 0.127M